OpenStudy (anonymous):

c(x)=250+6x+0.1x^2 marginal average cost at a production level of 30 units.

OpenStudy (amistre64):

i believe its the derivative of cost at 30 unites

OpenStudy (anonymous):

0+6+0.2x?

OpenStudy (amistre64):

very good, now when x=30, thats what .... 6 + 6 = 12 ?

OpenStudy (anonymous):

6+0.2(30) =12

OpenStudy (amistre64):

great! :)

OpenStudy (anonymous):

but the answer here is -0.18

OpenStudy (amistre64):

make sure your information that you posted is accurate

OpenStudy (anonymous):

it is.

OpenStudy (amistre64):

hmmm, then i missed something while reading. We are given the cost function C(x) the average cost function is therefore C(x)/x ... thats where i missed it the marginal average cost function is the derivative of C(x)/x .... not C(x)

OpenStudy (anonymous):

marginal cost at a pro lev. of 30 units.. =12 average cost func. is=250/x +6+0.1x marginal cost function is -250/x^2+0.1 ... i understand that part. except that marginal average cost

OpenStudy (anonymous):

im reviewing for my test. ;) hehe

OpenStudy (amistre64):

\[\frac{c(x)=250+6x+0.1x^2}{x}\]\[\frac{c(x)}{x}=250x^{-1}+6+0.1x\] \[D_x[\frac{c(x)}{x}]=-250x^{-2}+0+0.1\] \[-\frac{250}{30.30}+0.1\] \[-\frac{2.5}{9}+0.1\] \[-.27777....+0.1\] \[-.17777....\]

OpenStudy (amistre64):

the word marginal always brings to mind a derivative of what is being marginalized marginal (average cost) is the derivative of (average cost)

OpenStudy (anonymous):

how did u get 2.5/9? o.o

OpenStudy (anonymous):

ah nvm . i get it hehe

OpenStudy (anonymous):

ahhh!! ;)

OpenStudy (anonymous):

i got .377777777

OpenStudy (anonymous):

but if i subtract 0.1. i get .17777777778

OpenStudy (anonymous):

again. nvm. i just forgot to put - sign haha. thanks soo much! ;)